TWO BROTHERS HAVE A TOTAL AGE OF 3 YEARS.ONE IS 2 YEARS OLDER THAN THE OTHER.WHAT ARE THERE AGES
Answer by edjones(8007)

Let a be the age of the elder.
a+b=3
a-b=2
------add
2a=5
a=5/2= 2 1/2 yrs old
b= 6 months old
.
Ed